2a+6
------ = 6a+2
3

Answer by checkley77(12844)

2a+6
------ = 6a+2
3
2a+6=3(6a+2)
2a+6=18a+6
2a-16a=6-6
-14a-0
a=0/-14
a=0 ans:
Proof:
(2*0+6)/3=6*0+2
6/3=2
2=2